# : A Million-scale Real Prompt-Gallery Dataset for Text-to-Video Diffusion Models

### v1

@@ Abstract
The arrival of Sora marks a new era for text-to-video diffusion models, bringing significant advancements in video generation and potential applications. However, Sora, as well as other text-to-video diffusion models, highly relies on the prompts, and there is no publicly available dataset featuring a study of text-to-video prompts. In this paper, we introduce VidProM, the first large-scale dataset comprising 1.67 Million unique text-to-Video Prompts from real users. Additionally, the dataset includes 6.69 million videos generated by four state-of-the-art diffusion models and some related data. We initially demonstrate the curation of this large-scale dataset, which is a time-consuming and costly process. Subsequently, we show how the proposed VidProM differs from DiffusionDB, a large-scale prompt-gallery dataset for image generation. Based on the analysis of these prompts, we identify the necessity for a new prompt dataset specifically designed for text-to-video generation and gain insights into the preferences of real users when creating videos. Our large-scale and diverse dataset also inspires many exciting new research areas. For instance, to develop better, more efficient, and safer text-to-video diffusion models, we suggest exploring text-to-video prompt engineering, efficient video generation, and video copy detection for diffusion models. We make the collected dataset VidProM publicly available at GitHub and Hugging Face under the CC-BY-NC 4.0 License.
@@ 1 Introduction
The Sora [6] initializes a new era for text-to-video diffusion models, revolutionizing video generation with significant advancements. This breakthrough provides new possibilities for storytelling, immersive experiences, and content creation, as Sora [6] transforms textual descriptions into high-quality videos with ease. However, Sora [6] and other text-to-video diffusion models [9, 1, 3, 5] heavily relies on the prompts used. Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which hinders the development and evaluation of these models.
@@
In this paper, we present the first systematic research on the text-to-video prompts. Specifically, our efforts primarily focus on building the first text-to-video prompt-gallery dataset VidProM, conducting an in-depth comparison with DiffusionDB [22], analyzing prefered topics by users, and introducing new research directions based on our VidProM. The demonstration of VidProM is shown in Fig. 1.
@@
The first text-to-video prompt-gallery dataset. Our large-scale VidProM includes million unique text-to-video prompts from real users and million generated videos by state-of-the-art diffusion models. The prompts are from official Pika Discord channels, and the videos are generated by Pika [9], Text2Video-Zero [1], VideoCraft2 [3], and ModelScope [5]. We distribute the generation process across 10 servers, each equipped with 8 Nvidia V100 GPUs. Each prompt is embedded using the powerful text-embedding-3-large model from OpenAI and assigned six not-safe-for-work (NSFW) probabilities, consisting of toxicity, obscenity, identity attack, insult, threat, and sexual explicitness. We also add a Universally Unique Identifier (UUID) and a time stamp to each data point in our VidProM. In addition to the main dataset, we introduce a subset named VidProS, which consists of semantically unique prompts. In this subset, the cosine similarity between any two prompts is less than 0.8, ensuring a high level of semantic diversity.
@@
In-depth comparison with DiffusionDB and people preference analysis. We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [22]. By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [22] lies in: (1) Our VidProM contains more semantically unique prompts, which are embedded by a more advanced model and collected over a longer time span. (2) We collect videos through both web scraping and local generation, whereas DiffusionDB [22] only contains web-scraped images, resulting in our dataset being more time-consuming and costly to assemble. (3) The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer. These differences highlight the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models. Based on our analysis of the new prompt dataset, we conclude that (1) frequently occurring words include ‘modern’, ‘motion’, ‘close’, ‘forest’, and ‘sky’; and (2) preferred topics encompass humans, science fiction, and animals.
@@
Inspiring new research directions. The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up numerous exciting research directions. Researchers are focused on developing better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering. (2) For more efficient models, researchers can search for related prompts in our VidProM and reconstruct new videos from similar existing videos, thereby avoiding the need to generate videos from scratch. (3) For safer models, researchers can develop specialized models to distinguish generated videos from real videos to combat misinformation, and train video copy detection models to identify potential copyright issues. Apart from diffusion models, text-video pairs are also utilized in multimodal learning tasks such as video-text retrieval and video captioning. Our prompts and synthetic videos can help mitigate copyright concerns associated with using online videos and alleviate the difficulty of collecting high-quality video-text data.
@@
To sum up, this paper makes the following contributions: (1) We contribute the first text-to-video prompt-gallery dataset, VidProM, which includes million unique prompts from real users and million generated videos by state-of-the-art diffusion models. (2) We provide a detailed and in-depth comparison with the text-to-image prompt-gallery dataset, DiffusionDB, and highlight the necessity of VidProM as well as real users’ preference. (3) We reveal several exciting research directions inspired by VidProM and position it as a rich database for future studies.
@@ 3 Curating VidProM
Extracting and Embedding Prompts. The HTML files are then processed using regular expressions to extract prompts and time stamps. We subsequently filter out prompts used for image-to-video generation (because the images are not publicly available) and prompts without associated videos (these prompts may have been banned by Pika or hidden by the users). Finally, we remove duplicate prompts and assign a UUID to each one, resulting in a total of unique prompts. Because the text-to-video prompts are significantly complex and long, we use OpenAI’s text-embedding-3-large API, which supports up to tokens, to embed all of our prompts. We retain the original 3072-dimensional output of the text-embedding-3-large model, allowing users of our dataset to freely reduce their dimensionality according to their preferences using Principal Component Analysis (PCA).
@@
Scraping and Generating Videos. We enhance our VidProM diversity by not only scraping Pika videos from extracted links but also utilizing three state-of-the-art open-source text-to-video diffusion models for video generation. This process demands significant computational resources: we distribute the text-to-video generation across 10 servers, each equipped with 8 Nvidia V100 GPUs. It costs us approximately 50,631 GPU hours and results in 6.69 million videos (), totaling 24,414,747.8 seconds in duration. The breakdown of video lengths is as follows: 3.0 seconds for Pika, 2.0 seconds for Text2Video-Zero, 1.6 seconds for VideoCraft2, and 8.0 seconds for ModelScope.
Selecting Semantically Unique Prompts. Beyond general uniqueness, we introduce a new concept: semantically unique prompts. We define a dataset as containing only semantically unique prompts if, for any two arbitrary prompts, their cosine similarity calculated using text-embedding-3-large embeddings is less than . After semantic de-duplication, our VidProM still contains semantically unique prompts, and we denote it as VidProS. More semantically unique prompts imply covering a broader range of topics, increasing the diversity and richness of the content.
4 Comparison with DiffusionDB
In this section, we provide a comparison of the basic information between our VidProM and DiffusionDB [22], which is a large-scale prompt dataset for image generation. As shown in Table 1, the comparison is conducted in three aspects: prompts, images or videos, and licenses. We have the following observations:
@@ 4 Comparison with DiffusionDB
Licenses. Unlike DiffusionDB, which adheres to the CC0 1.0 license from the Stable Diffusion Discord, the prompts and videos generated by Pika in our VidProM are licensed under the CC BY-NC 4.0 license from the Pika Discord, allowing for non-commercial use with attribution. Additionally, similar to their original repositories, the videos from VideoCraft2, Text2Video-Zero, and ModelScope are released under the Apache license, the CreativeML Open RAIL-M license, and the CC BY-NC 4.0 license, respectively.
Aspects |
Details |
DiffusionDB [22] |
VidProM |
| Prompts |
No. of unique prompts |
No. of semantically unique prompts |
Embedding of prompts |
OpenAI-CLIP |
OpenAI-text-embedding-3-large |
Maximum length of prompts |
77 tokens |
8192 tokens |
Time span |
Aug 2022 |
Jul 2023 Feb 2024 |
| Images |
No. of images/videos |
million images |
million videos |
| or |
No. of sources |
| Videos |
Average repetition rate per source |
Collection method |
Web scraping |
Web scraping + Local generation |
GPU consumption |
V100 GPU hours |
Total seconds |
seconds |
Legal |
Open source license |
CC0 1.0 |
CC BY-NC 4.0 and others |
5 Prompt Analysis
In this section, we analyze the prompts in our VidProM dataset from two perspectives: (1) the necessity of introducing a new prompt dataset and (2) people’s preferences in video generation.
Why do we need a new prompt dataset for text-to-video generation? As there exists a prompt dataset (DiffusionDB) for text-to-image generation, we provide an in-depth analysis on how the prompts in our VidProM differ from those in DiffusionDB.
@@ 5 Prompt Analysis
Secondly, as shown in Fig. 4 (a), text-to-video prompts are generally more complex and longer than text-to-image prompts, due to the need to describe additional dimensions and dynamic changes. This phenomenon is also observed in the prompts used by Sora. For instance, the prompt for the ‘Tokyo Girl’ video contains words, while the longest prompt on the OpenAI official website comprises words. As depicted in Fig. 4 (b), our VidProM dataset prominently features this characteristic: (1) the number of prompts with more than words is nearly for our VidProM, compared to only about for DiffusionDB; and (2) our VidProM still has over prompts with more than words, whereas this number is close to for DiffusionDB.
Finally, as shown in Fig. 4 (c), the prompts from our VidProM dataset and DiffusionDB exhibit different distributions. We use the text-embedding-3-large model to re-extract the features of the prompts in DiffusionDB and randomly select prompt features from both our VidProM dataset and DiffusionDB, respectively. By visualizing with t-SNE [25], we find that these prompts have significantly different distributions and are nearly linearly separable.
What kind of videos are people interested in generating? We use two methods, i.e. WordCloud and WizMap [26], to analyze the text-to-video prompts preferred by people. WordCloud is a visual representation of text, in which the size of each word indicates its frequency of occurrence within the text. WizMap [26] is a visualization tool to navigate and interpret large-scale embedding spaces with ease. From the visualization results shown in Fig. 5, we conclude that:
(1) The popular words include modern, motion, close, forest, sky, and so on. In Fig. 5 (a), we can identify the most frequent words, which can be roughly divided into two groups: words describing the characteristics of the generated videos, indicating the desired style, such as ‘modern’, ‘motion’, and ‘close’; and words describing the subject of the generation, such as ‘forest’ and ‘sky’.
(2) The preferred topics include humans, science fiction, animals, and others. Firstly, the topic of humans is in the lower and middle parts of Fig. 5 (b): the middle part depicts general behaviors of men and women, such as walking, while the lower part showcases famous individuals like Joe Biden, Trump, and Jackson. Secondly, the science fiction topic is located in the upper part, including dragons flying in the sky from Chinese tradition and aliens with spaceships from the West. Thirdly, people also like cute animals, such as cats and dogs (in the lower right), as well as wild animals, such as tigers and lions (in the middle right). Finally, other topics, such as cars (on the right) and theology (in the lower left), are also mentioned in the prompts.
@@ 6 Inspiring New Research
Text-to-Video Diffusion Model Development aims to create diffusion models capable of converting textual descriptions into dynamic and realistic videos. The current methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] are trained on caption-(real)-video pairs. Two natural questions arise: (1) Will the domain gap between captions and prompts from real users hinder these models’ performance? (2) Can researchers train or distill new text-to-video diffusion models on prompt-(generated)-video pairs? The studies of training text-to-video diffusion models on our VidProM may provide answers to these two questions.
Text-to-Video Prompt Engineering is to optimize the interaction between humans and text-to-video models, ensuring that the models understand the task at hand and generate relevant, accurate, and coherent videos. The prompt engineering field has gained attention in large language models [52, 53], text-to-image diffusion models [54, 55], and visual in-context learning [56, 57]. However, as far as we know, there is no related research in the text-to-video community. Our VidProM provides an abundant resource for text-to-video prompt engineering.
@@
Fake Video Detection aims to distinguish between real videos and those generated by diffusion models. While there are some works [27, 28, 29, 30, 31, 32, 33, 34, 35, 36] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames. For instance, a model trained on images generated by Stable Diffusion [37] may fail to identify frames from videos generated by Pika [9] or Sora [6] as fake. (2) The efficiency problem: Currently, there is no detector that can take an entire video as input. As a result, we must use fake image detectors to examine all or representative frames, which can be time-consuming. With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from different diffusion models to further improve the models’ performance.
Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of their training sources. Videos generated with replicated content may infringe on the copyrights of the original videos. However, current video copy detection researches [38, 39, 40, 41] focus on detecting hand-crafted copies generated by transformations like horizontal flips and random crops, and overlook the challenge from diffusion models. Researchers can utilize our dataset VidProM as input queries and employ existing video copy detection models to pre-filter potential replications. Subsequently, these filtered videos can be manually annotated to determine whether they are indeed replications and regarded as training sources.

### v2

@@ Abstract
The arrival of Sora marks a new era for text-to-video diffusion models, bringing significant advancements in video generation and potential applications. However, Sora, along with other text-to-video diffusion models, is highly reliant on prompts, and there is no publicly available dataset that features a study of text-to-video prompts. In this paper, we introduce VidProM, the first large-scale dataset comprising 1.67 Million unique text-to-Video Prompts from real users. Additionally, this dataset includes 6.69 million videos generated by four state-of-the-art diffusion models, alongside some related data. We initially discuss the curation of this large-scale dataset, a process that is both time-consuming and costly. Subsequently, we underscore the need for a new prompt dataset specifically designed for text-to-video generation by illustrating how VidProM differs from DiffusionDB, a large-scale prompt-gallery dataset for image generation. Our extensive and diverse dataset also opens up many exciting new research areas. For instance, we suggest exploring text-to-video prompt engineering, efficient video generation, and video copy detection for diffusion models to develop better, more efficient, and safer models. The project (including the collected dataset VidProM and related code) is publicly available at https://vidprom.github.io under the CC-BY-NC 4.0 License.
@@ 1 Introduction
The Sora [1] initializes a new era for text-to-video diffusion models, revolutionizing video generation with significant advancements. This breakthrough provides new possibilities for storytelling, immersive experiences, and content creation, as Sora [1] transforms textual descriptions into high-quality videos with ease. However, Sora [1] and other text-to-video diffusion models [2, 3, 4, 5] heavily relies on the prompts used. Despite their importance, there is no publicly available dataset focusing on text-to-video prompts, which hinders the development of these models and related researches.
@@
In this paper, we present the first systematic research on the text-to-video prompts. Specifically, our efforts primarily focus on building the first text-to-video prompt-gallery dataset VidProM, analyzing the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models, and introducing new research directions based on our VidProM. The demonstration of VidProM is shown in Fig. 1.
@@
The first text-to-video prompt-gallery dataset. Our large-scale VidProM includes million unique text-to-video prompts from real users and million generated videos by state-of-the-art diffusion models. The prompts are from official Pika Discord channels, and the videos are generated by Pika [2], Text2Video-Zero [3], VideoCraft2 [4], and ModelScope [5]. We distribute the generation process across 10 servers, each equipped with 8 Nvidia V100 GPUs. Each prompt is embedded using the powerful text-embedding-3-large model from OpenAI and assigned six not-safe-for-work (NSFW) probabilities, consisting of toxicity, obscenity, identity attack, insult, threat, and sexual explicitness. We also add a Universally Unique Identifier (UUID) and a time stamp to each data point in our VidProM. In addition to the main dataset, we introduce a subset named VidProS, which consists of semantically unique prompts. That means, in this subset, the cosine similarity between any two prompts is less than 0.8, ensuring a high level of semantic diversity.
@@
The necessity of collecting a new prompt dataset specialized for text-to-video diffusion models. We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [6]. By analyzing the basic information and the prompts, we conclude that the differences between our VidProM and DiffusionDB [6] mainly lies in: (1) The semantics of our prompts are significantly different from those in DiffusionDB, with our text-to-video prompts generally being more dynamic, more complex, and longer. (2) Our VidProM contains more semantically unique prompts, which are embedded by a more advanced model and collected over a longer time span. (3) We collect videos through both web scraping and local generation, whereas DiffusionDB [6] only contains web-scraped images, resulting in our dataset being more time-consuming and costly to assemble.
@@
Inspiring new research directions. The introduction of our new text-to-video prompt-gallery dataset, VidProM, opens up many exciting research directions. With the help of our VidProM, researchers can develop better, more efficient, and safer text-to-video diffusion models: (1) For better models, researchers can utilize our VidProM as a comprehensive set of prompts to evaluate their trained models, distill new models using our prompt-(generated)-video pairs, and engage in prompt engineering. (2) For more efficient models, researchers can search for related prompts in our VidProM and reconstruct new videos from similar existing videos, thereby avoiding the need to generate videos from scratch. (3) For safer models, researchers can develop specialized models to distinguish generated videos from real videos to combat misinformation, and train video copy detection models to identify potential copyright issues. Apart from diffusion models, text-video pairs can also be utilized in multimodal learning tasks such as video-text retrieval and video captioning. Our prompts and synthetic videos can help mitigate copyright concerns associated with using online videos and alleviate the difficulty of collecting video-text data.
@@
To sum up, this paper makes the following contributions: (1) We contribute the first text-to-video prompt-gallery dataset, VidProM, which includes million unique prompts from real users and million generated videos by state-of-the-art diffusion models. (2) We highlight the necessity of collecting a new prompt dataset specialized for text-to-video diffusion models by comparing our VidProM with DiffusionDB. (3) We reveal several exciting research directions inspired by VidProM and position it as a rich database for future studies.
@@ 3 Curating VidProM
Extracting and Embedding Prompts. The HTML files are then processed using regular expressions to extract prompts and time stamps. We subsequently filter out prompts used for image-to-video generation (because the images are not publicly available) and prompts without associated videos (these prompts may have been banned by Pika or hidden by the users). Finally, we remove duplicate prompts and assign a UUID to each prompt, resulting in a total of unique prompts. Because the text-to-video prompts are significantly complex and long, we use OpenAI’s text-embedding-3-large API, which supports up to tokens, to embed all of our prompts. We retain the original 3072-dimensional output of the text-embedding-3-large model, allowing people to freely reduce their dimensionality according to their preferences using Principal Component Analysis (PCA).
@@
Scraping and Generating Videos. We enhance our VidProM diversity by not only scraping Pika videos from extracted links but also utilizing three state-of-the-art open-source text-to-video diffusion models for video generation. This process demands significant computational resources: we distribute the text-to-video generation across 10 servers, each equipped with 8 Nvidia V100 GPUs. It costs us approximately 50,631 GPU hours and results in 6.69 million videos (), totaling 14,381,289.8 seconds in duration. The breakdown of video lengths is as follows: 3.0 seconds for Pika, 1.6 seconds for VideoCraft2, 2.0 seconds for Text2Video-Zero, and 2.0 seconds for ModelScope.
Selecting Semantically Unique Prompts. Beyond general uniqueness, we introduce a new concept: semantically unique prompts. We define a dataset as containing only semantically unique prompts if, for any two arbitrary prompts, their cosine similarity calculated using text-embedding-3-large embeddings is less than . After semantic de-duplication (see a detailed description of this process in the Appendix (Section A)), our VidProM still contains semantically unique prompts, and we denote it as VidProS. More semantically unique prompts imply covering a broader range of topics, increasing the diversity and richness of the content.
| Aspects | Details | DiffusionDB [6] | VidProM |
| Prompts | No. of unique prompts | ||
| No. of semantically unique prompts | |||
| Embedding of prompts | OpenAI-CLIP | OpenAI-text-embedding-3-large | |
| Maximum length of prompts | 77 tokens | 8192 tokens | |
| Time span | Aug 2022 | Jul 2023 Feb 2024 | |
| Images | No. of images/videos | million images | million videos |
| or | No. of sources | ||
| Videos | Average repetition rate per source | ||
| Collection method | Web scraping | Web scraping + Local generation | |
| GPU consumption | - | V100 GPU hours | |
| Total seconds | - | seconds |
4 The Necessity of Introducing VidProM
We notice that there exists a text-to-image prompt-gallery dataset, DiffusionDB [6]. This section highlights how our VidProM is different from this dataset from two aspects, i.e. basic information and semantics of prompts.
4.1 Basic Information
In Table 1, we provide a comparison of the basic information between our VidProM and DiffusionDB [6]. We have the following observations:
@@ 4.1 Basic Information
4.2 Semantics of prompts
In this section, we analyze how the semantics of prompts in our VidProM dataset are different from DiffusionDB [6].
@@ 4.2 Semantics of prompts
Secondly, as shown in Fig. 4 (a) and (b), text-to-video prompts are generally more complex and longer than text-to-image prompts, due to the need to describe additional dimensions and dynamic changes. This phenomenon is also observed in the prompts used by Sora. For instance, the prompt for the ‘Tokyo Girl’ video contains words, while the longest prompt on the OpenAI official website comprises words. Our VidProM dataset prominently features this characteristic: (1) the number of prompts with more than words is nearly for our VidProM, compared to only about for DiffusionDB; and (2) our VidProM still has over prompts with more than words, whereas this number is close to for DiffusionDB.
Finally, as shown in Fig. 4 (c) and Fig. 5, the prompts from our VidProM dataset and DiffusionDB exhibit different distributions. We use the text-embedding-3-large model to re-extract the features of the prompts in DiffusionDB. For visualizing with t-SNE [25], we randomly select prompt features from both our VidProM dataset and DiffusionDB, respectively. We find that these prompts have significantly different distributions and are nearly linearly separable. Beyond the traditional t-SNE [25], we use a recent method, WizMap [26], to analyze the topics preferred by people using all prompts from our VidProM and DiffusionDB. WizMap [26] is a visualization tool to navigate and interpret large-scale embedding spaces with ease. From the visualization results shown in Fig. 5, we conclude that, despite some overlaps, there are distinct differences in their interests. For example, both groups are interested in topics like cars, food, and cute animals. However, text-to-image users tend to focus on generating art paintings, whereas text-to-video users show little interest in this area. Instead, text-to-video users are more inclined to create general human activities, such as walking, which are rarely produced by text-to-image users.
@@ 5 Inspiring New Research
Text-to-Video Diffusion Model Development aims to create diffusion models capable of converting textual descriptions into dynamic and realistic videos. The current methods [3, 7, 4, 8, 5, 1, 9, 10, 2, 11, 12] are trained on caption-(real)-video pairs. Two natural questions arise: (1) Will the domain gap between captions and prompts from real users (see the evidence of existing domain gap in the Appendix (Section B)) hinder these models’ performance? For instance, of the training data for Open-Sora-Plan v1.0.0 [31] consist of landscape videos, leading to the trained model’s suboptimal performance in generating scenes involving humans and cute animals. (2) Can researchers train or distill new text-to-video diffusion models on prompt-(generated)-video pairs? Although some studies indicate that this approach may lead to irreversible defects [32], as we will likely run out of data on websites and some methods are proposed to prevent such collapse [33], exploring synthetic data remains a valuable endeavor. The studies of training text-to-video diffusion models with our VidProM may provide answers to these two questions.
Text-to-Video Prompt Engineering is to optimize the interaction between humans and text-to-video models, ensuring that the models understand the task at hand and generate relevant, accurate, and coherent videos. The prompt engineering field has gained attention in large language models [34, 35], text-to-image diffusion models [36, 37], and visual in-context learning [38, 39]. However, as far as we know, there is no related research in the text-to-video community. Our VidProM provides an abundant resource for text-to-video prompt engineering. In the Appendix (Section C), we train a large language model on our VidProM for automatic text-to-video prompt completion.
@@
Fake Video Detection aims to distinguish between real videos and those generated by diffusion models. While there are some works [40, 41, 42, 43, 44, 45, 46, 47, 48, 49] focusing on fake image detection, fake video detection presents unique challenges: (1) The generalization problem: Existing fake image detectors may not generalize well to video frames (see experiments in the Appendix (Section D)). For instance, a model trained on images generated by Stable Diffusion [50] may fail to identify frames from videos generated by Pika [2] or Sora [1] as fake. (2) The efficiency problem: Currently, there is no detector that can take an entire video as input. As a result, to achieve higher accuracy, we may use fake image detectors to examine all or representative frames, which can be time-consuming. With our VidProM, researchers can (1) train specialized Fake Video Detection models on millions of generated videos, and (2) use millions of prompts to generate more videos from more diffusion models to further improve the detection performance.
Video Copy Detection for Diffusion Models aims to answer whether videos generated by diffusion models replicate the contents of existing ones (see a detailed discussion in the Appendix (Section E)). Videos generated with replicated content may infringe on the copyrights of the original videos. However, current video copy detection researches [51, 52, 53, 54] focus on detecting hand-crafted copies generated by transformations like horizontal flips and random crops, and overlook the challenge from diffusion models. Researchers can utilize our dataset VidProM as input queries and employ existing video copy detection models to pre-filter potential replications. Subsequently, these filtered videos can be manually annotated to determine whether they are indeed replications and regarded as training sources.